OpenStudy (anonymous):

can you help my son solve this one 2/3x-1=5/3x+1

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[\frac{2}{3x-1}=\frac{5}{3x+1}\] best bet it to cross multiply and get \[2(3x+1)=5(3x-1)\]

OpenStudy (anonymous):

good from there?

OpenStudy (anonymous):

\[2(3x+1)=5(3x-1)\] \[6x+2=15x-5\] \[2+5=15x-6x\] \[7=9x\] \[\frac{7}{9}=x\]

OpenStudy (anonymous):

satellite73, you're a genius. now I've seen it written down, a light has come on in my head. That's his homework done!